OpenStudy (anonymous):

You are given the sides of a triangle and need to determine if the triangle is acute, obtuse or right. 8. Sides have lengths 9 in., 7 in., and 8 in. 9. Sides have lengths 36 ft., 39 ft., and 15 ft. 10. Sides have lengths 9 m, 33 m, and 41 m. Could someone please explain how to do this. I don't want answers, I want to know how to do this.

OpenStudy (anonymous):

I think question 8 wants works like this. a^2 + b^2 = c^2 6^2 +7^2 = 8^2 36 + 49 = 64 But 36 + 49 does not equal 64, it equals 85. So I'm left with C = 64 but A + B = 85. Does that make this an obtuse or an acute triangle?

OpenStudy (anonymous):